A baseket contains 6blue,2 red,4 greenand 3 yellow balls
1. if 2 balls are picked at random, what is the probability that either both are
green or both are yellow?
2. If 5 balls are picked at random, what is the probability that at least one is blue.
3. If 2 balls are picked at random, what is the probability that both are blue?
4. If 4 balls are picked at random what is the probability that 2 are red and 2 are green
5.if 3 balls are picked at random what is the probability that none is yellow

Answer by stanbon(75887)

A baseket contains 6 blue,2 red,4 green and 3 yellow balls
1. if 2 balls are picked at random, what is the probability that either both are
green or both are yellow?
---
P(2green OR 2yellow) = P(2green) + P(2yellow)
= [4C2/15C2]+[3C2/15C2] = [6+3]/105 = 9/105
-----------------------------------------------
2. If 5 balls are picked at random, what is the probability that at least one is blue.
P(at least one blue) = 1-P(none blue)
= 1 - [9C5/15C5]
= 1 - 0.0420
= 0.9580
---------------------------------------------------
3. If 2 balls are picked at random, what is the probability that both are blue?
P(2blue) = [6C2/15C2]
-------------------------------
4. If 4 balls are picked at random what is the probability that 2 are red and 2 are green
P(2red AND 2green) = P(2green|2red)*P(2red) = [3C2/13C2][2C2/15C2]
--------------------------------------
5.if 3 balls are picked at random what is the probability that none is yellow
P(no yellow) = 12C3/15C3
==============================
Cheers,
Stan H.
===============